OpenStudy (anonymous):

Write the equation of the line through each point. Use slope intercept form. (1, -1); parallel to y = 2/5x – 3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

None:/

OpenStudy (skullpatrol):

Do you know what "slope intercept" form looks like?

OpenStudy (anonymous):

I know the formula its y-y1 = m (x – x1) but i'm not sure how to plug it into this

OpenStudy (skullpatrol):

$$\huge y-y_1 = m (x – x_1)$$ $$\huge y=mx+b$$ $$\huge y = (2/5)x – 3$$ $$\huge x_1=1$$ $$\huge y_1=-1$$

OpenStudy (anonymous):

Alright and then what do we do?

OpenStudy (skullpatrol):

What is the value of m?

OpenStudy (anonymous):

I'm awful at math sorry:/ How do we find m?

OpenStudy (skullpatrol):

Look at the second and third equations and match the symbol m to the number.

OpenStudy (skullpatrol):